10. (2x2 + 8xy + y2) dx + (2x2 + xy3/3) dy = 0

determine which equations are exact and solve them
Answer #1

1o. (2x² + 8 xy + y 2) du + (2x2 4 xy²/3) dy zo. M dut Ndly a equation is exact if 3 Boy - 2.0+ B3.04 + 2y = 831+2y 2N = 4x +